Status Report

SOHO Daily Meeting Minutes for Wednesday, 03 July 2002

By SpaceRef Editor
July 3, 2002
Filed under , ,

* RHESSI target for today is AR 0017; target for tomorrow
is TBD.

* Friday’s SVM reserved time has been scheduled for 13:30 to 15 UT
instead of 14-15 UT. As usual no NRT is available during this time.
* Security scans for EOF and EAF scheduled for 3 July at 10am

* Thursday 4 July is a holiday. There will be a student ECS
operator in the office from about 11am-4pm.

* CONTOUR launched today

Spacecraft Status: Nominal

Spacecraft Anomalies: None

Accomplished Activities: RSL, VIRGO, SWAN, New Tracking Star

Planned Activities: VIRGO

Upcoming Operations: None

Ground Anomalies: None

07/02 DOY 183 – 07/03 DOY 184

EIT daily On-Line Solar Status and observations:

LASCO daily On-Line Solar Status and observations:

Web Page for Planning:

CDS: Nominal. Flare study and JOP136 on AR0017. Tomorrow:
sunspot study on AR0019. Friday: EUV AR intensity
study on AR0019.

UVCS: Nominal. Coronal hole/streamer interface study and
mini-synoptic for today and Friday. Tomorrow: 24-hour

LASCO: Nominal. C2/C3 synoptics, JOP159 today through Friday.
Starting Friday for approximately a month, LASCO will
be taking an hourly pB C2 image. One C2 image, one C3
image, and a 195 image will be lost each hour for this.

EIT: Nominal. 195 CME watch and synoptic.

MDI: Nominal. Full disk magnetograms and dopplergrams.

TRACE: Nominal. Supporting DOT/THEMIS observations on AR0019
for today through Friday. There will not be a local
TRACE planner until next Wednesday, so if anyone needs
to contact TRACE, send email or call Dick Shine at

SpaceRef staff editor.